Consider the line 3x+8y=5 What is the slope of a line parallel to this line? What is the slope of a line perpendicular to this line?
I am so confused
Tiff

Answer by Alan3354(69443)

Consider the line 3x+8y=5 What is the slope of a line parallel to this line? What is the slope of a line perpendicular to this line?
--------------
Put 3x+8y=5 into slope-intercept form.
That means solve for y.
--------
3x+8y=5
y = (-3/8)x + 5/8
----
The slope is -3/8
All parallel lines have the same slope, -3/8
---------
The slope of lines perpendicular is the negative inverse, = +8/3
------
That's all there is to that.